OpenStudy (anonymous):

1/8+3/4 a.1/2 b.1/3 c.1/4 d.2/3

OpenStudy (anonymous):

(3/4)*(2/2)= 6/8 1/8+6/8 = 7/8 so i dont get how those are the only choices